OpenStudy (shadowcain):

Martina walked 3/4 of a mile in 3/5 of an hour. At this rate, how far can Martina walk in one hour?

OpenStudy (freckles):

\[\frac{\frac{3}{4} \text{ mile }}{\frac{3}{5} \text{ hour }} =\frac{ ? \text{ mile }}{1 \text{ hour }}\] find the ? mark

OpenStudy (shadowcain):

A. 1 2/3 miles B. 9/20 of a mile C. 4/5 of a mile D. 1 1/4 miles

OpenStudy (freckles):

let me know if you need anymore help

OpenStudy (shadowcain):

I keep coming up with 11/4, but that's not one of my options

OpenStudy (freckles):

did you try to solve the equation I gave you above?

OpenStudy (freckles):

\[\frac{\frac{3}{4}}{\frac{3}{5}}=\frac{x}{1} \\ \text{ or } \frac{\frac{3}{4}}{\frac{3}{5}}=x\] you are just dividing 3/4 by 3/5

OpenStudy (freckles):

can you try finishing from here? \[x=\frac{\frac{3}{4}}{\frac{3}{5}} \cdot \frac{\frac{5}{3}}{\frac{5}{3}}\] let me know what you get after trying to do this multiplication on top and bottom

OpenStudy (shadowcain):

I'm not sure how to do that

OpenStudy (freckles):

\[x=\frac{\frac{3}{4} \cdot \frac{5}{3}}{\frac{3}{5} \cdot \frac{5}{3}} \\ x=\frac{\frac{3(5)}{4(3)}}{\frac{3(5)}{5(3)}}\]

OpenStudy (freckles):

hint 15/15 is just one anything divided by 1 is that anything \[x=\frac{3(5)}{4(3)}\]

OpenStudy (shadowcain):

15/12 = 1 1/4 miles

OpenStudy (freckles):

yep

OpenStudy (shadowcain):

Thanks

OpenStudy (freckles):

np